OpenStudy (milo123):

Is this 66? The height of a cone-shaped container is 18 centimeters and its radius is 7 centimeters. Meg fills the container completely with honey. Every day Meg uses 21 cubic centimeters of honey from the container to make honey cookies. After how many days will the container be completely empty? Round your answer to the nearest whole number. (Use π = 3.14.) 44 66 78 92

OpenStudy (milo123):

@Elsa213

OpenStudy (studygurl14):

First find the total volume of the container.

OpenStudy (studygurl14):

Once you find that, use this formula V - 21x = 0 V = volume Solve for x, the number of days

OpenStudy (milo123):

so i'm taking that as a redo it it's not 66 lol?

OpenStudy (will.h):

I'll check it for you

OpenStudy (studygurl14):

No, it's not 66

OpenStudy (studygurl14):

Volume of a cone: \(\huge\frac{1}{3}\Large r^2h\pi\)

OpenStudy (milo123):

923.62?

OpenStudy (will.h):

V = Pi* r^2* h/3 V= pi * 7^2 * 18/3 V = 923.63

OpenStudy (will.h):

And then as StudyGurl mentioned use V - 21x = 0

OpenStudy (studygurl14):

Now divide it by 21 to find how many days it will take

OpenStudy (will.h):

923.63/21 = 44

OpenStudy (milo123):

44

OpenStudy (studygurl14):

yep :)

OpenStudy (will.h):

Good

OpenStudy (milo123):

Thanks guys!!

OpenStudy (studygurl14):

anytime :)

OpenStudy (will.h):

Yw